OpenStudy (awesome781):

@timo86m Help me study!

OpenStudy (anonymous):

what is the q?

OpenStudy (anonymous):

make a system of linear inequalities :)

OpenStudy (awesome781):

How?

OpenStudy (anonymous):

x+y+z=5000 z=900 y>z x>y that is your system :)

OpenStudy (awesome781):

aaaahhhhhT_T

OpenStudy (anonymous):

i am not 100% sure but you have to solve for x :)

OpenStudy (awesome781):

So if I plug z in it would be x+y+900=5000 Nd I would have to find x nd y?

OpenStudy (anonymous):

lets back track a bit you know how i got x+y+z=5000 z=900 y>z x>y RIGHT?

OpenStudy (awesome781):

Yea.....

OpenStudy (anonymous):

then lets reduce it more by plugging in x+y=4100 <- we simplified and took out 900 y>900 x>y

OpenStudy (awesome781):

Ok.... I'm following....

OpenStudy (anonymous):

the q askes what is the least number of votes that candidate x could have received?

OpenStudy (awesome781):

So we need to subtract?

OpenStudy (anonymous):

yes i am thinking tho x>y>900 wee need to solve for x+y=4100 :) we pic y as 901 and now we solve x+901=4100 Think about it z had 900 and the next highest number is 901 so we need to use that. Cuzz half votes or quarter votes dont count :)

OpenStudy (anonymous):

but we not done yet tho :)

OpenStudy (anonymous):

cuzz that would be the max However we can use the max to solve for min. Maybe my way took an extra step but it got us there.

OpenStudy (awesome781):

U_U ok.

OpenStudy (anonymous):

x+901=4100 x=3199 would be max :) but they want min

OpenStudy (awesome781):

So we have to convert it .....to minutes right?

OpenStudy (anonymous):

to to minimum :)

OpenStudy (anonymous):

hold on gotta go get a pizza

OpenStudy (awesome781):

Ok......I'm waiting ......hurry...lol^_^

OpenStudy (awesome781):

Jk...........take ur time:)

OpenStudy (anonymous):

not to butt in, but can't you spit it in half and add 1

OpenStudy (awesome781):

@satellite73 what do u mean?

OpenStudy (anonymous):

you have 900 votes accounted for leaving 4100 you want the LEAST amount for the highest vote getter

OpenStudy (anonymous):

half of 4100 is 2050 so if they both got 2050 they would have the same total but if one had 2049 and the other had 2051 then that is the smallest the winner could have

OpenStudy (anonymous):

i.e. if X had less than 2051 then it would not be more than candidate Y

OpenStudy (awesome781):

Whatt I understand.......ur saying is to divide each and subtract one?

OpenStudy (anonymous):

x>y>900 x+y=4100, the smallest number x can be is 2051, satellite is right

OpenStudy (anonymous):

so you have 4100 and you need to divide it so that the largest part is as small as possible: at 2051<2049 x (which is 2051) is as small as it can be while still satisfying the inequality.

OpenStudy (anonymous):

I mean 2051>2049

OpenStudy (awesome781):

Jus making sure......u divide by 2 from 4100, right?

OpenStudy (anonymous):

The possibilities go from 4099>1 to 4098>2 to 4097>3 ... and so on, the object being to count all the way down until the leading number is as small as possible

OpenStudy (awesome781):

How did u get the smallest number?

OpenStudy (anonymous):

if you keep counting from where I left off you will eventually get to 2051>2049, and if you go past that the inequality no longer works (2050>2050 is false because 2050 is not greater than 2050, they are equal)

OpenStudy (awesome781):

Ok thnx for your explanation:)

OpenStudy (anonymous):

You're welcome.

OpenStudy (awesome781):

So the answer is 2051

OpenStudy (anonymous):

yep